# Costless correction of chain based nested sampling parameter estimation in gravitational wave data and beyond

### v1

@@ 6 Conclusions
In nested sampling, the evidence error bar is dominated by the unknown exact prior volumes of the ‘shells’ between iso-likelihood contours. This error affects parameter estimation too, but here there is an additional source of error, from using the parameter value of a single sample, $f(\theta_{i})$, as a proxy for the average parameter value over the entire iso-likelihood contour defined by that sample, $\tilde{f}(X_{i})$ (Chopin & Robert, 2010). Though the prior volume error is well understood and accurately estimated from the ‘simulated weights method’, the added stochasticity due to parameter variation over a given contour is typically ignored, despite the fact that this can be a significant source of error in the analysis of any dataset, including gravitational wave signals (Thrane & Talbot, 2019).
Here we have proposed two novel approaches to account for this additional error using the extra likelihood calls made at run-time by any chain based sampler. Though these ‘phantom points’ are not suitable for use in evidence estimation, they are valuable in parameter inference for exploring the variation in a parameter value over individual contours. This is the first demonstration of the use of nested sampling phantom points for inference. We have also shown how to use phantom points to verify the accuracy of the error bars on a given parameter, by splitting the set of the phantom points in two and checking for convergence in the resulting estimates. This could be a useful approach at runtime in tuning the chain length of chain based samplers to speed up nested sampling whilst still ensuring independent samples. In other MCMC procedures, such as Hamiltonian Monte Carlo (Carpenter et al., 2017; Cabezas et al., 2024), the emphasis is placed on tracking convergence on specific parameters of interest, and dynamic nested sampling (Higson et al., 2018b; Speagle, 2020) provides a way to apply this in nested sampling too.
Furthermore, we have shown how this error impacts estimates of credible intervals and coverage plots, such as the $p$-$p$ plot commonly used in gravitational wave analyses. Though this source of error is often secondary to the statistical error due to the finite number of events sampled for a typical coverage plot, the two become comparable in size for larger numbers of analysed events. This highlights the importance of accounting for the additional nested sampling parameter estimation error when interpreting coverage plots.

### v2

@@ 5 Conclusions
In nested sampling, the evidence error bar is dominated by the unknown prior volumes of the ‘shells’ between iso-likelihood contours. This affects parameter estimation too, but here there is an additional source of uncertainty, from using the parameter value of a single sample, $f(\theta_{i})$, as a proxy for the average parameter value over the entire iso-likelihood contour defined by that sample, $\tilde{f}(X_{i})$ (Chopin & Robert, 2010). Though the prior volume uncertainty is well understood and accurately estimated from the ‘simulated weights method’, the added stochasticity due to parameter variation over a given contour is typically ignored, despite the fact that this can be a significant source of uncertainty in the analysis of any dataset, including gravitational wave signals (Thrane & Talbot, 2019).
Here we have proposed two novel approaches to account for this additional uncertainty using the extra likelihood calls made at run-time by any chain based sampler. Though these ‘phantom points’ are not suitable for use in evidence estimation, they are valuable in parameter inference for exploring the variation in a parameter value over individual contours. This is the first demonstration of the use of nested sampling phantom points for inference. We have also shown how to use phantom points to verify the accuracy of the error bars on a given parameter, by splitting the set of the phantom points in two and checking for convergence in the resulting estimates. This could be a useful approach at runtime in tuning the chain length of chain based samplers to speed up nested sampling whilst still ensuring independent samples. In other MCMC procedures, such as Hamiltonian Monte Carlo (Carpenter et al., 2017; Cabezas et al., 2024), the emphasis is placed on tracking convergence on specific parameters of interest, and dynamic nested sampling (Higson et al., 2018b; Speagle, 2020) provides a way to apply this in nested sampling too.
Furthermore, we have shown how this uncertainty impacts estimates of credible intervals and coverage plots, such as the $p$-$p$ plot commonly used in gravitational wave analyses. Though this source of uncertainty is often secondary to the statistical uncertainty due to the finite number of events sampled for a typical coverage plot, the two become comparable in size for larger numbers of analysed events. This highlights the importance of accounting for the additional nested sampling parameter estimation uncertainty when interpreting such plots.
@@ Acknowledgements
Appendix A Other parameters
As well as the chirp mass and luminosity distance, the methods presented in this paper were used to estimate uncertainties on other parameters. Since both methods perform similarly well to each other on all the parameters, for conciseness we present the plots for the likelihood binning method only.
As discussed above, on some parameters neither method works significantly better than the standard methods, and this is largely driven by the lack of sufficient uncorrelated phantom points in these parameters. The mass ratio (Figure 13) is one such parameter. This can be rectified by running a longer chain during sampling. As discussed in Section 4.3, we can verify that the error bar on the mass ratio parameter computed from a single run with the default chain length is not correct by following the method described in Section 3.3.3. Performing a run with a chain length of $100$ instead, Figures 14 and 15 corroborate that the original chain length of $20$ was too short to produce sufficiently many uncorrelated phantom points for use in these methods, and a much longer chain must be used.
However, on other parameters such as the total mass, both methods successfully calculate the correct error bars, as they did with the chirp mass (Figure 16). Crucially, the total mass is not one of the four sampled parameters in the run, and thus demonstrates the robustness of these methods to arbitrary functions of the sampling parameters.
Figure 13: The mass ratio is an example of a parameter on which our methods (purple) did not significantly outperform standard methods (red) of uncertainty estimation for the default chain length settings we used. As with the luminosity distance, this can be rectified by setting the chain length to be longer.
Figure 14: For a run with a chain length of $100$, the parameter estimates using the likelihood binning method from the first 2020 points in the chain do not match those from the last $80$. This indicates that the first 2020 points are perhaps too correlated in their mass ratio values to use in these methods.
Figure 15: Running a longer chain corrects the issue. We can safely use the phantoms points from position $50$ onwards in the chain, and our methods will give the correct bars.
Figure 16: The total mass is a function of two sampled parameters, the chirp mass and the mass ratio. Our methods (purple) are able to successfully capture the true uncertainty on such a parameter, whereas the simulated weights method (red) is not.
Appendix B Other samplers and settings
The methods presented in this paper apply to any chain-based sampler, but the results may be dependent on the exact sampler settings used. Below, we use PolyChord with the default settings for slice sampling in dynesty (Speagle, 2020) to demonstrate that these settings need to be carefully decided using the metrics presented in this paper in order to obtain valid uncertainty estimates.
For the slice sampler setting in dynesty, the default chain length is $3+\textrm{ndims}$, in contrast to PolyChord’s $5\times\textrm{ndims}$. We note here that dynesty does not default to slice sampling until the number of dimensions is above 20. However, the default sampler for 4 dimensions is not a chain-based one and thus our methods would not apply. Additionally, many users may choose correctly to change that default as rejection sampling can become inefficient in much lower dimensions too. dynesty also doesn’t have a clustering algorithm as part of its nested sampler, and so to emulate this code, we run PolyChord with a chain length of $7$ for our $4$ parameter problem and turn off clustering in the settings. The corresponding results are presented in Figures 1720.
Figure 17: The equivalent of Figure 3 for the new settings. The middle left panel shows the estimated error bars on the evidence using the ‘simulated weights’ method and the bottom left shows a histogram of the percentiles in which the ‘true evidence’, estimated over 100 runs, lies. The K-S test $p$-value for a uniform distribution is $0.0621$. On the right hand side panels, the equivalent figures for the chirp mass estimates are shown. These show qualitatively the same results as for the longer chain.
Figure 18: The likelihood binning method applied to these phantom points results in a distribution that is still not uniform, though the error bars are wider. This is due to the chain length not being long enough to produce sufficiently many uncorrelated phantom points.
Figure 19: The phantom runs method also does not significantly improve the uncertainty estimate on the chirp mass, and here performs slightly worse than likelihood binning. This is again due to the chain length being too short.
Neither the likelihood binning nor the phantom runs method works particularly well here, and this is because although the chain length may be long enough to produce unbiased estimates of the evidence, there aren’t sufficiently many uncorrelated phantom points to use. We can verify this by following the method discussed in Section 3.3.3, comparing the estimates obtained from the first 4 phantom points in every chain to those from the last 3. The resulting distributions have a K-S 2-sample test statistic of $0.31$, with a $p$-value of $2.2\times 10^{-4}$, indicating that they may not be in agreement. This is a sign that a longer chain needs to be run.
Figure 20: In order to test whether the chain length setting of $3+\textrm{ndims}$, the default for slice sampling in dynesty, is long enough to apply these methods successfully, we can follow the methods presented in Section 3.3.3. Splitting the phantom points into two sets, according to their position in the chains, we compare the estimates produced by each set independently. In this case, the K-S 2-sample test statistic between the resulting distributions was $0.31$, with a $p$-value of $2.2\times 10^{-4}$.
The two methods in this paper should work for any chain-based sampler, but this example shows the importance of the chain length setting to the validity of the results. In order to get accurate parameter estimation uncertainties, it may be necessary to run a longer chain than is needed for unbiased evidence estimates. Again, the metrics presented in Section 3.3.3 will enable users to manually set the appropriate chain length.